Automobile repair costs continue to rise with the average cost now at $367 per repair (U.S. News & World Report website, January 5, 2015). Assume that the cost for an automobile repair is normally distributed with a standard deviation of $88. Answer the following questions about the cost of automobile repairs. a) What is the probability that the cost will be more than $450? b) What is the probability that the cost will be less than $250? c) What is the probability that the cost will be between $250 and $450? d) If the cost for your car repair is in the lower 5% of automobile repair charges. what is vour cost?
Answer #1

Mean (H)-367 and Standarddeviation (o)-88 a) The probability thatthe cost will be more than S450 We have to find P (X> 450) When x= 450, Z (x-H)_(450-367 83_ 0.9432 Therefore, P(X> 450)-P (Z> 0.9432) = P (05 Z < o)-P (0 = 0.5-0.3264 Z 0.9432) [From Normal table] P(X> 450) 0.1736 b) The probability thatthe cost will be less than S250 We have to find P(X< 250) When x- 250, 250-367 117--1.3295 Z-- Therefore, P(X<250)-P (Z<1.3295) 0.5-0.4082 [From Normal table] P(X< 250) 0.0918c) The probability that the cost will be between S250 and S450 We have to find P (250 SX S 450) P (250 SX 450) P (-13295 S Zs